Brandtly Industries invests a large sum of money in R&D; as a result, it retains and reinvests all of its earnings. In other words, Brandtly does not pay any dividends, and it has no plans to pay dividends in the near future. A major pension fund is interested in purchasing Brandtly's stock. The pension fund manager has estimated Brandtly's free cash flows for the next 4 years as follows: $3 million, $6 million, $9 million, and $15 million. After the fourth year, free cash flow is projected to grow at a constant 3%. Brandtly's WACC is 15%, the market value of its debt and preferred stock totals $80 million; the firm has $14 million in non-operating assets; and it has 24 million shares of common stock outstanding.

  1. What is the present value of the free cash flows projected during the next 4 years? Do not round intermediate calculations. Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.
    $  

  2. What is the firm's horizon, or continuing, value? Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.
    $  

  3. What is the market value of the company's operations? Do not round intermediate calculations. Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.
    $  

    What is the firm's total market value today? Do not round intermediate calculations. Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.
    $  

  4. What is an estimate of Brandtly's price per share? Do not round intermediate calculations. Round your answer to the nearest cent.
    $  
Answer #1

a.Present value of cash flows = 3,000,000/(1.15) + 6,000,000/(1.15)2 + 9,000,000/(1.15)3 + 15,000,000/(1.15)4

= $21,639,502.43

b.Horizon Value = Free cash flow in Year 5/(WACC – growth rate)

= 15,000,000(1.03)/(15%-3%)

= $128,750,000

c.Market Value is equal to present value of all future free cash flows

= 21,639,502.43 + 128,750,000/(1.15)4

= $95,252,732.80

d.Value of firm = 95,252,733

Add: Non-operating assets = 14,000,000

Less: value of debt = $80,000,000

Value of Equity = $29,252,733

Number of shares = 24,000,000

Price per share = $1.2189